# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 2.2. Koch–Tataru spaces
We state two equivalent definitions of the space $BMO^{-1}(\mathbb{R}^{d})$. Then the periodic version $BMO^{-1}(\mathbb{T}^{d})$ is defined as the space of distributions on $\mathbb{T}^{d}$ whose periodic extensions are in $BMO^{-1}(\mathbb{T}^{d})$, with the norm defined as the norm of the extension.
@@
where the infimum is taken over cubes $Q$ in $\mathbb{R}^{d}$. An equivalent definition of the $BMO^{-1}$ norm is
@@
$\|f\|_{\dot{F}_{1,2}^{1}(\mathbb{T}^{d})}$ $=\left\|\left(\sum_{N}N^{2}|P_{N}f|^{2}\right)^{2}\right\|_{L^{1}(\mathbb{T}^{d})}.$
@@
The second formulation of $BMO^{-1}$ motivates the definition of the natural path space for the heat equation from such data, first introduced by Koch and Tataru [KochTataru2001]. For $T\in(0,\infty]$, let
$X_{T}(\mathbb{R}^{d})\coloneqq\left\{u\in L_{loc}^{2}(\mathbb{R}^{d}\times[0,T];\mathbb{R}^{d}):\operatorname{div}u=0,\,\|u\|_{X_{T}}<\infty\right\}$
@@ 2.3. Weak solutions
The notion of a weak solution provides the minimal analytic framework in which the Navier–Stokes equations can be interpreted for arbitrary finite-energy initial data. Since the pioneering work of Leray [Leray1934], it has been known that weak solutions exist globally in time for any $d\geq 2$.
@@ 3.1. Frequency scales
Define parameters $A>1$, which will be taken as large as needed throughout the proofs, and $b>1$, which can be safely assigned as $b=3$. Fix also an $m_{*}\in\mathbb{N}$ to be specified later. From here we define the frequency scales as follows: fix the minimum frequency $N_{1,0}=1$ at $(j,k)=(1,0)$, while for all other $k\in\mathbb{N}$ and $j\in\mathcal{J}_{d}$, define
$N_{j,k}=\left\{\begin{split}&\big\lceil A^{b^{k+(j-1)/J_{d}}}\big\rceil,\quad&d=2,\\ &\big\lceil A^{b^{k}}\big\rceil,\quad&d\geq 3,\end{split}\right$. (3.1)
@@
$M_{j,k}=\left\{\begin{split}&\big\lceil A^{\gamma b^{k+(j-1)/J_{d}}}\big\rceil,\quad&d=2,\\ &\big\lceil A^{\gamma b^{k}}\big\rceil,\quad&d\geq 3,\end{split}\right$.
@@
$\gamma>b^{-1/J_{d}}$
and taking $A$ still larger as needed, we arrange that
@@ 3.2. Geometry of the building blocks
Choose some $\theta_{j},\eta_{j}\in\mathbb{S}^{d-1}\cap\mathbb{Q}^{d}$ such that $\theta_{j}\cdot\eta_{j}=0$ for all $j\in\mathcal{J}_{d}$ and $k\geq 0$, and $(\theta_{j}\otimes\theta_{j})_{j\in\mathcal{J}_{d}}$ are linearly independent and generate $\operatorname{Id}$ with positive coefficients. Fixing $m_{*}$ in Section 3.1 to be such that $m_{*}\eta_{j}\in\mathbb{Z}^{d}$ for all $j$, we ensure that $N_{j,k}\eta_{j}\in\mathbb{Z}^{d}$ for all $j$ and $k$.
For $\rho<1/10$, define $\mathcal{C}_{j,k}(\rho)$ the $2\pi/M_{j,k}$-periodic cylinder of radius $\rho M_{j,k}^{-1}$ whose axis is the periodic line $\mathbb{R}\theta_{j}\pmod{2\pi\mathbb{Z}^{d}/M_{j,k}}$. Fix $\delta_{0}>0$ small enough (depending on $d$ but not $k$) so that
@@
The proof is similar to [CoiculescuPalasek2025]*Lemma 3.3. Let $\rho_{k}\in C^{\infty}(\mathbb{T}^{d};[0,1])$ be $2\pi/M_{1,k}$-periodic with
$\rho_{k}\equiv 1\quad\text{on }\mathcal{C}_{1,k}(3\delta_{0})$
@@
$\operatorname{supp}\rho_{k}\subset\mathcal{C}_{1,k}(4\delta_{0}).$
@@
$\|\rho_{k}\|_{L^{1}}\leq|\mathcal{C}_{1,k}(4\delta_{0})|\leq 1/10.$
By improved Hölder’s inequality (see, e.g., Lemma 2.1 in [ModenaSz18]),
@@ 3.3. The velocity potential
To avoid issues related to non-local operators in the construction, we define the principal part of the velocity field as the Laplacian of a particular potential.
Let $\phi_{k}$ be a standard mollifier at length scale $\ell_{k}$ where $N_{1,k+1}^{-1}\ll\ell_{k}\ll N_{J_{d},k}^{-1}$. Define the sequence of profiles
$\psi_{j,0}(x)$ $\coloneqq N_{j,0}^{-2}a_{j,0}(x)\theta_{j}\sin(N_{j,0}\eta_{j}\cdot x)$ (3.8)
@@
$\sum_{j}a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}=2\mathcal{D}\sum_{j}N_{j,k}\psi_{j,k}+p\operatorname{Id}$ (3.10)
@@
$\|\nabla^{n}\psi_{j,k}\|_{L^{\infty}}\lesssim_{n}N_{j,k}^{-2+n}$ (3.13)
@@
$S=\sum_{j\in\mathcal{J}_{d}}\Gamma_{j}^{2}(S)\theta_{j}\otimes\theta_{j}$
for all $S\in\mathrm{Sym}^{d}(\mathbb{R})$ in a $c_{0}$-neighborhood of $\operatorname{Id}$, where $\Gamma_{j}:B(\operatorname{Id},c_{0})\to\mathbb{R}$ are smooth. For $k=0$, we define $a_{1,0}(x)=1$, and $a_{j,0}(x)=0$ for $j\neq 1$.
Recursively, we define
@@
and $\psi_{j,k}$ is defined in (3.9) in terms of $a_{j,k}$. The absolute constant $c$ will be chosen small enough so that
$c\|S_{k}\|_{\infty}\leq c_{0}$ (3.14)
@@
$\|a_{j,k}\|_{\infty}\lesssim c^{-\frac{1}{2}}.$
for all $j$ and $k$. We can also get rough estimates on the derivatives of $\psi_{j,k}$ by shifting all the derivative onto the mollifier:
@@
$\|\nabla^{n}S_{k}\|_{\infty}\lesssim\sum_{j}N_{j,k}\|\nabla^{n+1}\psi_{j,k}\|_{\infty}\lesssim N^{-1}_{J_{d},k}\ell_{k}^{-n-1}c^{-\frac{1}{2}},$
In particular, since $N^{-1}_{J_{d},k}\ell^{-1}_{k}=o(k)$,
$c\|S_{k}\|_{\infty}\lesssim c^{\frac{1}{2}}N^{-1}_{J_{d},k}\ell_{k}^{-1}\lesssim c^{\frac{1}{2}}.$
Hence condition (3.14) is satisfied for $c$ small enough and the rank-one decomposition can be applied.
Before we obtain optimal bounds on the derivatives, we continue with rough estimates. Using Leibniz rule we obtain
$\|\nabla^{n}_{x}\Gamma_{j}(\operatorname{Id}+cS_{k}(x))\|_{\infty}$ $\lesssim_{n}\sum_{i=1}^{n}\sum_{\alpha_{1}+\cdots+\alpha_{i}=n}c^{i}\|\nabla^{\alpha_{1}}S_{k}\|_{\infty}\cdots\|\nabla^{\alpha_{i}}S_{k}\|_{\infty}\|\nabla_{i}\Gamma_{j}\|_{L^{\infty}}$
$\lesssim c^{\frac{1}{2}}N^{-1}_{J_{d},k}\ell_{k}^{-n-1}.$
Since $M_{J_{d},k}\leq N_{J_{d},k}\leq\ell_{k}^{-1}$,
$\begin{split}\|\nabla^{n}a_{j,k+1}\|_{\infty}&=\big\|\nabla^{n}\Big(c^{-\frac{1}{2}}\chi_{k+1}(x)\Gamma_{j}\big(\operatorname{Id}+cS_{k}(x)\big)\Big)\big\|_{\infty}\\ &\lesssim_{n}c^{-\frac{1}{2}}\|\nabla^{n}_{x}\Gamma_{j}(\operatorname{Id}+cS_{k}(x))\|_{\infty}\\ &\lesssim_{n}N^{-1}_{J_{d},k}\ell_{k}^{-n-1}.\end{split}$
Now we are in a position to obtain optimal bounds on the derivatives of $S_{k}$. Since $N^{-1}_{J_{d},k-1}\leq\ell_{k-1}^{-1}\leq M_{j,k}\leq N_{j,k}$,
$\begin{split}\|\nabla^{n}\psi_{j,k}\|_{\infty}&\lesssim_{n}N_{j,k}^{-2}\sum_{i=0}^{n}\|\nabla^{i}(a_{j,k}\varphi_{j,k})\|_{\infty}N_{j,k}^{n-i}\\ &\lesssim_{n}c^{-\frac{1}{2}}N_{j,k}^{n-2},\end{split}$
which in turn improves bounds on $S_{k}$
$\begin{split}\|\nabla^{n}S_{k}\|_{\infty}&\lesssim\sum_{j}N_{j,k}\|\nabla^{n+1}\psi_{j,k}\|_{\infty}\\ &\lesssim c^{-\frac{1}{2}}N_{J_{d},k}^{n},\end{split}$
@@
$\operatorname{supp}\psi_{j,k}=(\operatorname{supp}{a_{j,k}}\cap\operatorname{supp}\varphi_{j,k})+B(0,\ell_{k}),$
@@ 4.1. Error bounds on the velocity
Then for any $\varepsilon_{0}>0$, $\alpha\in(0,\frac{1}{10})$, and $\bar{n}\in\mathbb{N}$, the following bounds hold for all $t\geq 0$
$\|\nabla^{n}(\bar{R}_{k}(t)-\mathcal{I}_{k}(t))\|_{L^{\infty}}\leq\varepsilon_{0}N^{-\alpha}_{1,k+1}(t^{-\frac{n}{2}+\alpha}+1),\qquad t\geq 0,\quad n=1,2,\dots,\bar{n},$
$\|\bar{R}_{k}(t)-\mathcal{I}_{k}(t)\|_{L^{\infty}}\lesssim_{\varepsilon}\varepsilon_{0}N_{1,k+1}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}}N_{J_{d},k}^{\varepsilon},\qquad\text{if}\qquad d=2,$
$\|\bar{R}_{k}(t)-\mathcal{I}_{k}(t)\|_{L^{\infty}}\lesssim\varepsilon_{0}N_{1,k+1}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}},\qquad\text{if}\qquad d\geq 3,$
provided $A$, $b$, and $m$ are large enough.
@@
for some sufficiently small $\varepsilon>0$. In the case $d\geq 3$ the above bound can be improved. Indeed, in this case $N_{j,k}=N_{1,k}$ for all $j$. Then the summation in $j$ becomes, by (3.10),
$\begin{split}\nabla^{n}\mathbb{Q}\sum_{j}(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j})&=\nabla^{n}\mathcal{R}\mathbb{P}\operatorname{div}\left(2\mathcal{D}\sum_{j}N_{1,k}\psi_{j,k}+p\operatorname{Id}\right)\\ &=\nabla^{n}\mathcal{R}\Delta\mathbb{P}\sum_{j}N_{1,k}\psi_{j,k}\\ &=\nabla^{n}\mathcal{R}\Delta\mathbb{P}\sum_{j}N_{1,k}^{-1}\phi_{k}*(a_{j,k}(x)\varphi_{j,k}(x)\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x))\\ &=\nabla^{n}\mathcal{R}\Delta\sum_{j}N_{1,k}^{-1}\phi_{k}*(a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x))\\ &\quad-\nabla^{n}\mathcal{R}\nabla\sum_{j}N_{1,k}^{-1}\phi_{k}*(\nabla(a_{j,k}\varphi_{j,k})\cdot\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x)),\end{split}$
@@
$\begin{split}\|\nabla^{n}(e^{t\Delta}-\operatorname{Id})\mathbb{Q}(a^{2}_{j,k}\theta_{j}\otimes\theta_{j})\|_{\infty}&\lesssim t\|\nabla^{n}\Delta\mathbb{Q}(a^{2}_{j,k}\theta_{j}\otimes\theta_{j})\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}tN_{J_{d},k-1}^{n+2+\varepsilon}.\end{split}$
@@
$\begin{split}\|\nabla^{n}(\mathcal{I}_{k}^{<t_{k}}-\tilde{\mathcal{I}}_{k}^{<t_{k}})\|_{\infty}&=\Big\|e^{(t-t_{k})\Delta}\int_{0}^{t_{k}}e^{(t_{k}-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)\,ds\\ &\quad-e^{(t-t_{k})\Delta}\int_{0}^{t_{k}}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)\,ds\Big\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}\int_{0}^{t_{k}}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}t_{k}N_{J_{d},k}^{n+2+\varepsilon}\,ds\\ &\lesssim t_{k}N_{J_{d},k}^{n+2+\varepsilon}(1-e^{-2N_{j,k+1}^{2}t_{k}})\\ &\lesssim t_{k}N_{J_{d},k}^{n+2+\varepsilon},\end{split}$
@@
$\frac{1}{2}\sum_{j}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)$ $=\sum_{j}N_{j,k}\nabla\odot\mathbb{P}\psi_{j,k}$
$=\sum_{j}N_{j,k}(\tilde{\mathcal{D}}\psi_{j,k}-(\operatorname{Id}-2\Delta^{-1}\nabla\otimes\nabla)\operatorname{div}\psi_{j,k}),$ (4.4)
@@
$\begin{split}\|\nabla^{n}\tilde{\mathcal{I}}_{k}^{1}\|_{L^{\infty}}&\lesssim\sum_{j}N_{j,k}^{n}e^{-N_{j,k}^{2}t}(e^{N_{j,k}^{2}t_{k}}-1)\lesssim\sum_{j}N_{j,k}^{n+2}t_{k}e^{-N_{j,k}^{2}t}\end{split}$ (4.5)
@@
$\|\nabla^{n}\tilde{\mathcal{I}}_{k}^{2}\|_{\infty}$ $\lesssim\sum_{j}N_{j,k}^{n}\left(M_{j,k}N_{j,k}^{-1}e^{-N_{j,k}^{2}t/4}+N_{j,k}^{-m}M_{j,k}^{m}\right)$

### v2

@@ 3.2. Koch–Tataru spaces
We state two equivalent definitions of the space $BMO^{-1}(\mathbb{R}^{d})$. Then the periodic version $BMO^{-1}(\mathbb{T}^{d})$ is defined as the space of distributions on $\mathbb{T}^{d}$ whose periodic extensions are in $BMO^{-1}(\mathbb{R}^{d})$, with the norm defined as the norm of the extension.
@@
where the supremum is taken over cubes $Q$ in $\mathbb{R}^{d}$. An equivalent definition of the $BMO^{-1}$ norm is
@@
$\|f\|_{\dot{F}_{1,2}^{1}(\mathbb{T}^{d})}$ $=\left\|\left(\sum_{N}N^{2}|P_{N}f|^{2}\right)^{\frac{1}{2}}\right\|_{L^{1}(\mathbb{T}^{d})}.$
@@
The second formulation of $BMO^{-1}$ motivates the definition of the natural path space for the heat equation from such data, first introduced by Koch and Tataru [32]. For $T\in(0,\infty]$, let
$X_{T}(\mathbb{R}^{d})\coloneqq\left\{u\in L_{loc}^{2}(\mathbb{R}^{d}\times[0,T];\mathbb{R}^{d}):\operatorname{div}u=0,\,\|u\|_{X_{T}}<\infty\right\},$
@@ 3.3. Weak solutions
The notion of a weak solution provides the minimal analytic framework in which the Navier–Stokes equations can be interpreted for arbitrary finite-energy initial data. Since the pioneering work of Leray [36], it has been known that weak solutions exist globally in time for any $d\geq 2$.
@@
Note that weak solutions also satisfy the mild formulation, i.e,
$u=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-s)\Delta}\mathbb{P}\operatorname{div}(u\otimes u)(s)\,ds,$
in the sense of distributions, [18].
@@ 4.1. Frequency scales
We introduce the parameters $b>1$, to be chosen large depending only on $d\geq 2$, and $A>1$, to be chosen large depending on all other parameters. Fix also an $m_{*}\in\mathbb{N}$ to be specified later. From here we define the frequency scales as follows: fix the minimum frequency $N_{1,0}=1$ at $(j,k)=(1,0)$, while for all other $k\in\mathbb{N}$ and $j\in\mathcal{J}_{d}$, define
$N_{j,k}=\left\{\begin{aligned} &m_{*}\big\lceil A^{b^{k+(j-1)/J_{d}}}\big\rceil,\quad&d=2,\\ &m_{*}\big\lceil A^{b^{k}}\big\rceil,\quad&d\geq 3,\end{aligned}\right$. (4.1)
@@
$M_{j,k}=\left\{\begin{aligned} &\big\lceil A^{\gamma b^{k}}\big\rceil,\quad&j=1\text{ or }d\geq 3,\\ &\big\lceil A^{\gamma b^{(j-1)/J_{d}}}\big\rceil M_{1,k},\quad&j\geq 2\text{ and }d=2,\end{aligned}\right$. (4.2)
@@
$\gamma>b^{-1/J_{d}},$
and taking $A$ still larger as needed, we arrange that for $d=2$,
@@ 4.2. Geometry of the building blocks
Choose some $\theta_{j},\eta_{j}\in\mathbb{S}^{d-1}\cap\mathbb{Q}^{d}$ such that $\theta_{j}\cdot\eta_{j}=0$ for all $j\in\mathcal{J}_{d}$, and $(\theta_{j}\otimes\theta_{j})_{j\in\mathcal{J}_{d}}$ are linearly independent and generate $\operatorname{Id}$ with positive coefficients. Fixing $m_{*}$ in Section 4.1 to be such that $m_{*}\eta_{j}\in\mathbb{Z}^{d}$ for all $j$, we ensure that $N_{j,k}\eta_{j}\in\mathbb{Z}^{d}$ for all $j$ and $k$.
For $\rho<1/10$, define $\mathcal{C}_{j,k}(\rho)$ to be the $2\pi/M_{j,k}$-periodic cylinder of radius $\rho M_{j,k}^{-1}$ whose axis is the periodic line $\mathbb{R}\theta_{j}\pmod{2\pi\mathbb{Z}^{d}/M_{j,k}}$. Fix $\delta_{0}>0$ small enough (depending on $d$ but not $k$) so that
@@
The proof is similar to [12]*Lemma 3.3. Recall that, by definition, $\mathcal{C}_{j,k}(r)$ is a $2\pi/M_{j,k}$-periodic cylinder. Since $M_{1,k}$ divides $M_{j,k}$ for all $j\in\mathcal{J}_{d}$ (by (4.2)), the cylinders are all in fact $2\pi/M_{1,k}$-periodic. Thus we may arrange cutoff functions $\rho_{k}\in C^{\infty}(\mathbb{T}^{d};[0,1])$ that are $2\pi/M_{1,k}$-periodic with
$\rho_{k}\equiv 1\quad\text{on }\bigcup_{j\in\mathcal{J}_{d}}\mathcal{C}_{j,k}(3\delta_{0}),$
@@
$\operatorname{supp}\rho_{k}\subset\bigcup_{j\in\mathcal{J}_{d}}\mathcal{C}_{j,k}(4\delta_{0}).$
@@
$\|\rho_{k}\|_{L^{1}}\leq\sum_{j\in\mathcal{J}_{d}}|\mathcal{C}_{j,k}(4\delta_{0})|\leq 1/10.$
By improved Hölder’s inequality (see, e.g., Lemma 2.1 in [40]),
@@ 4.3. The velocity potential
To avoid issues related to nonlocal operators in the construction, we define the principal part of the velocity field as the Laplacian of a particular potential.
Let $\phi_{k}$ be a standard mollifier at length scale $\ell_{k}:=N_{1,k}^{-\frac{1}{2}}N_{1,k+1}^{-\frac{1}{2}}$. We will need $\ell_{k}\leq N_{J_{d},k}^{-1}$, which clearly holds provided $b$ is large enough. Define the sequence of profiles
$\psi_{j,0}(x)$ $\coloneqq N_{j,0}^{-2}a_{j,0}(x)\theta_{j}\sin(N_{j,0}\eta_{j}\cdot x),$ (4.10)
@@
$\sum_{j}a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}=2\mathcal{D}\sum_{j}N_{j,k}\psi_{j,k}+p\operatorname{Id},$ (4.12)
@@
$\|\nabla^{n}\psi_{j,k}\|_{L^{\infty}}\lesssim_{n}N_{j,k}^{-2+n},$ (4.15)
@@
$S=\sum_{j\in\mathcal{J}_{d}}\Gamma_{j}^{2}(S)\theta_{j}\otimes\theta_{j},$
for all $S\in\mathrm{Sym}^{d}(\mathbb{R})$ in a $c_{0}$-neighborhood of $\operatorname{Id}$, where $\Gamma_{j}:B(\operatorname{Id},c_{0})\to\mathbb{R}$ are smooth.
We define $a_{1,0}(x)=1$, $a_{j,0}(x)=0$ for $j\neq 1$, and, recursively, for $k\geq 0$,
@@
and $\psi_{j,k}$ is defined in (4.10)–(4.11) in terms of $a_{j,k}$. The absolute constant $c$ will be chosen small enough so that
$c\|S_{k}\|_{\infty}\leq c_{0},$ (4.16)
@@
$\|a_{j,k}\|_{\infty}\lesssim c^{-\frac{1}{2}},$
for all $j$ and $k$. We can also get rough estimates on the derivatives of $\psi_{j,k}$ by shifting all the derivatives onto the mollifier:
@@
$\|\nabla^{n}S_{k}\|_{\infty}\lesssim\sum_{j}N_{j,k}\|\nabla^{n+1}\psi_{j,k}\|_{\infty}\lesssim N^{-1}_{1,k}\ell_{k}^{-n-1}c^{-\frac{1}{2}}.$
Before we obtain optimal bounds on the derivatives, we continue with rough estimates. Using the Faà di Bruno formula,
$\|\nabla^{n}_{x}\Gamma_{j}(\operatorname{Id}+cS_{k}(x))\|_{\infty}$ $\lesssim_{n}\sum_{i=0}^{n}\sum_{\alpha_{1}+\cdots+\alpha_{i}=n}c^{i}\|\nabla^{\alpha_{1}}S_{k}\|_{\infty}\cdots\|\nabla^{\alpha_{i}}S_{k}\|_{\infty}\|\nabla^{i}\Gamma_{j}\|_{L^{\infty}}$
$\lesssim\sum_{i=0}^{n}\sum_{\alpha_{1}+\cdots+\alpha_{i}=n}c^{i}(N_{1,k}^{-1}\ell_{k}^{-\alpha_{1}-1}c^{-\frac{1}{2}})\cdots(N_{1,k}^{-1}\ell_{k}^{-\alpha_{i}-1}c^{-\frac{1}{2}})$
$\lesssim\sum_{i=0}^{n}c^{\frac{i}{2}}N_{1,k}^{-i}\ell_{k}^{-n-i}$
$\lesssim c^{\frac{n}{2}}N^{-n}_{1,k}\ell_{k}^{-2n}.$
Since $M_{J_{d},k}\leq N_{J_{d},k}\leq\ell_{k}^{-1}$, we may neglect the terms in which derivatives fall on $\chi_{k+1}$, which yields
$\begin{split}\|\nabla^{n}a_{j,k+1}\|_{\infty}&=\big\|\nabla^{n}\Big(c^{-\frac{1}{2}}\chi_{k+1}(x)\Gamma_{j}\big(\operatorname{Id}+cS_{k}(x)\big)\Big)\big\|_{\infty}\\ &\lesssim_{n}c^{-\frac{1}{2}}\|\nabla^{n}_{x}\Gamma_{j}(\operatorname{Id}+cS_{k}(x))\|_{\infty}\\ &\lesssim_{n}c^{\frac{n-1}{2}}N^{-n}_{1,k}\ell_{k}^{-2n}.\end{split}$
Now we are in a position to obtain optimal bounds on the derivatives of $S_{k}$. Since $\ell_{k-1}^{-1}=N_{1,k-1}^{\frac{1}{2}}N_{1,k}^{\frac{1}{2}}$ and $M_{j,k}\leq N_{j,k}$,
$\begin{split}\|\nabla^{n}\psi_{j,k}\|_{\infty}&\lesssim_{n}N_{j,k}^{-2}\sum_{i=0}^{n}\|\nabla^{i}(a_{j,k}\varphi_{j,k})\|_{\infty}N_{j,k}^{n-i}\\ &\lesssim_{n}N_{j,k}^{-2}\sum_{i=0}^{n}(c^{\frac{i-1}{2}}N^{-i}_{1,k-1}\ell_{k-1}^{-2i}+c^{-\frac{1}{2}}M_{j,k}^{i})N_{j,k}^{n-i}\\ &\lesssim_{n}N_{j,k}^{-2}c^{-\frac{1}{2}}N_{j,k}^{n},\end{split}$
which in turn improves the estimate of $S_{k}$:
$\begin{split}\|\nabla^{n}S_{k}\|_{\infty}&\lesssim\sum_{j}N_{j,k}\|\nabla^{n+1}\psi_{j,k}\|_{\infty}\\ &\lesssim\sum_{j}N_{j,k}N_{j,k}^{-2}c^{-\frac{1}{2}}N_{j,k}^{n+1}\\ &\lesssim\sum_{j}c^{-\frac{1}{2}}N_{j,k}^{n}\\ &\lesssim c^{-\frac{1}{2}}N_{J_{d},k}^{n},\end{split}$
@@
$\operatorname{supp}\psi_{j,k}\subset(\operatorname{supp}{a_{j,k}}\cap\operatorname{supp}\varphi_{j,k})+B(0,\ell_{k}),$
@@ 5.1. Error bounds on the velocity
Then for any $\varepsilon_{0}>0$, $\alpha\in(0,\frac{1}{10})$, and $\bar{n}\in\mathbb{N}$, the following bounds hold for all $t\geq 0$:
$\|\nabla^{n}(\bar{R}_{k}(t)-\mathcal{I}_{k}(t))\|_{L^{\infty}}\leq\varepsilon_{0}N^{-\alpha}_{1,k}(t^{-\frac{n}{2}+\alpha}+1),\qquad n=1,2,\dots,\bar{n},$
$\|\bar{R}_{k}(t)-\mathcal{I}_{k}(t)\|_{L^{\infty}}\lesssim_{\varepsilon}\varepsilon_{0}N_{1,k}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}}N_{J_{d},k}^{\varepsilon},\qquad\text{if}\qquad d=2,$
$\|\bar{R}_{k}(t)-\mathcal{I}_{k}(t)\|_{L^{\infty}}\lesssim\varepsilon_{0}N_{1,k}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}},\qquad\text{if}\qquad d\geq 3,$
provided $A$ and $b$ are large enough.
@@
for any $\varepsilon>0$. In the case $d\geq 3$ the above bound can be improved. Indeed, in this case $N_{j,k}=N_{1,k}$ for all $j$. Then the summation in $j$ becomes, by (4.12),
$\begin{split}\nabla^{n}\mathbb{Q}\sum_{j}(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j})&=\nabla^{n}\mathcal{R}\mathbb{P}\operatorname{div}\left(2\mathcal{D}\sum_{j}N_{1,k}\psi_{j,k}+p\operatorname{Id}\right)\\ &=2\nabla^{n}\mathcal{R}\Delta\mathbb{P}\sum_{j}N_{1,k}\psi_{j,k}\\ &=2\nabla^{n}\mathcal{R}\Delta\mathbb{P}\sum_{j}N_{1,k}^{-1}\phi_{k}*(a_{j,k}(x)\varphi_{j,k}(x)\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x))\\ &=2\nabla^{n}\mathcal{R}\Delta\sum_{j}N_{1,k}^{-1}\phi_{k}*(a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x))\\ &\quad-2\nabla^{n}\mathcal{R}\nabla\sum_{j}N_{1,k}^{-1}\phi_{k}*(\nabla(a_{j,k}\varphi_{j,k})\cdot\theta_{j}\sin(N_{1,k}\eta_{j}\cdot x)),\end{split}$
@@
$\begin{split}\|\nabla^{n}(e^{t\Delta}-\operatorname{Id})\mathbb{Q}(a^{2}_{j,k+1}\theta_{j}\otimes\theta_{j})\|_{\infty}&\lesssim t\|\nabla^{n}\Delta\mathbb{Q}(a^{2}_{j,k+1}\theta_{j}\otimes\theta_{j})\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}tN_{J_{d},k}^{n+2+\varepsilon}.\end{split}$
@@
$\begin{split}\|\nabla^{n}(\mathcal{I}_{k}^{<t_{k}}-\tilde{\mathcal{I}}_{k}^{<t_{k}})\|_{\infty}&=\Big\|e^{(t-t_{k})\Delta}\int_{0}^{t_{k}}e^{(t_{k}-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)\,ds\\ &\quad-e^{(t-t_{k})\Delta}\int_{0}^{t_{k}}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)\,ds\Big\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}\int_{0}^{t_{k}}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}t_{k}N_{J_{d},k}^{n+2+\varepsilon}\,ds\\ &\lesssim t_{k}N_{J_{d},k}^{n+2+\varepsilon},\end{split}$
@@
$\frac{1}{2}\sum_{j}\mathbb{Q}\left(a_{j,k+1}^{2}(x)\theta_{j}\otimes\theta_{j}\right)$ $=2\sum_{j}N_{j,k}\nabla\odot\mathbb{P}\psi_{j,k}$
$=\sum_{j}N_{j,k}(\tilde{\mathcal{D}}\psi_{j,k}+(\operatorname{Id}-2\Delta^{-1}\nabla\otimes\nabla)\operatorname{div}\psi_{j,k}),$ (5.4)
@@
$\begin{split}\|\nabla^{n}\tilde{\mathcal{I}}_{k}^{1}\|_{L^{\infty}}&\lesssim\sum_{j}N_{j,k}^{n}e^{-N_{j,k}^{2}t}(e^{N_{j,k}^{2}t_{k}}-1)\lesssim\sum_{j}N_{j,k}^{n+2}t_{k}e^{-N_{j,k}^{2}t},\end{split}$ (5.5)
@@
$\|\nabla^{n}\tilde{\mathcal{I}}_{k}^{2}\|_{\infty}$ $\lesssim\sum_{j}N_{j,k}^{n}\left(M_{j,k}N_{j,k}^{-1}e^{-N_{j,k}^{2}t/4}+N_{j,k}^{-m}M_{j,k}^{m}\right),$